# Plain Transformers Can be Powerful Graph Learners

### v2

@@ Abstract
Transformers have attained outstanding performance across various modalities, owing to their simple but powerful scaled-dot-product (SDP) attention mechanisms. Researchers have attempted to migrate Transformers to graph learning, but most advanced Graph Transformers (GTs) have strayed far from plain Transformers, exhibiting major architectural differences either by integrating message-passing or incorporating sophisticated attention mechanisms. These divergences hinder the easy adoption of training advances for Transformers developed in other domains. Contrary to previous GTs, this work demonstrates that the plain Transformer architecture can be a powerful graph learner. To achieve this, we propose to incorporate three simple, minimal, and easy-to-implement modifications to the plain Transformer architecture to construct our Powerful Plain Graph Transformers (PPGT): (1) simplified $L_{2}$ attention for measuring the magnitude closeness among tokens; (2) adaptive root-mean-square normalization to preserve token magnitude information; and (3) a simple MLP-based stem for graph positional encoding. Consistent with its theoretical expressivity, PPGT demonstrates noteworthy realized expressivity on the empirical graph expressivity benchmark, comparing favorably to more complicated competitors such as subgraph GNNs and higher-order GNNs. Its outstanding empirical performance across various graph datasets also justifies the practical effectiveness of PPGT.
@@ 1 Introduction
Transformers have achieved excellent performance across various domains, from language Vaswani et al., (2017); Devlin et al., (2019); Brown et al., (2020) to vision Dosovitskiy et al., (2021); Touvron et al., 2021a ; Touvron et al., 2021b , and are well known for their reduced dependency on inductive bias as well as stronger flexibility and generalizability Dosovitskiy et al., (2021). The recent success of transformer-based, multi-modal, large language models (LLMs) OpenAI, (2024); Dubey et al., (2024) has also brought the once unattainable dream of building artificial general intelligence closer to reality. However, these modalities are primarily data on Euclidean spaces, while non-Euclidean domains are still under-explored. Non-Euclidean domains like graphs are important for describing physical systems with different symmetry properties, from protein structure prediction Jumper et al., (2021) to complex physics simulation Sanchez-Gonzalez et al., (2018). Therefore, many researchers have attempted to migrate Transformer architectures for graph learning for the future integration into multimodal foundation models.
Nevertheless, unlike other Euclidean domains, the characteristics of graphs make the naive migration challenging. Following the suboptimal performance of an early attempt Dwivedi and Bresson, (2021), researchers introduced significant architectural modifications to make Transformers perform well for graph learning. These include: implicit/explicit message-passing mechanisms Kreuzer et al., (2021); Rampášek et al., (2022), edge-updating Kim et al., (2022); Hussain et al., (2022); Ma et al., (2023), and sophisticated non-SDP (scaled-dot-product) attention designs Chen et al., (2022); Ma et al., (2023). These complexities hinder the integration of advances for Transformers arising from other domains and impede progress towards potential future unification of multi-modalities.
In this work, we step back and rethink the difficulties that a plain Transformer architecture faces when processing graph data. Instead of adding significant architectural modifications, we propose three minimalist but effective modifications to empower plain Transformers with advanced capabilities for learning on graphs (as shown in Fig. 1(a)). This results in our proposed Powerful Plain Graph Transformers (PPGTs), which have remarkable empirical and theoretical expressivity for distinguishing graph structures and superior empirical performance on real-world graph datasets. These modifications can be incorporated without significantly changing the architecture of the plain Transformer. Thus, our proposal both retains the simplicity and generality of the Transformer and offers the potential of facilitating cross-modality unification.
@@ 2.1 Graph Learning
[…]
@@
The philosophy of Graph Transformers (GTs) is to learn graph representations using positional encodings (PEs) rather than operating directly on the original input graphs. Zhang et al., 2023b provide a theoretical framework (GD-WL) that demonstrates that GTs’ stronger expressivity stems from finer-grained information beyond 1-WL encoded in graph PEs. Zhang et al., 2024b demonstrate that GD-WL not only applies to distance-like or relative graph PEs but also enables analysis of absolute PEs (e.g., Laplacian PE) by transforming them into their relative counterparts. Concurrently, Ma et al., (2024) interpret graph PE as pseudo-coordinates in graph spaces that mimic manifolds, demonstrating that it can extend beyond graph distance and is not restricted to the Transformer architecture (e.g., convolutions on pseudo-coordinates). We follow the same philosophy and show that with a simple stem on pseudo-coordinates, even plain Transformers can achieve outstanding expressive power in graph learning.
@@ 2.2 Limitations in Plain Transformer Architectures
LayerNorm (LN Ba et al., (2016)) and Root-Mean-Square-Norm (RMSN Zhang and Sennrich, (2019)) are two widely used token-wise normalization techniques in Transformer-based models that effectively control token magnitudes:
$\text{LN}(\mathbf{x})=\frac{\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}\|}\odot\boldsymbol{\gamma}+\boldsymbol{\beta}\,,\quad$ $\text{RMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\odot\boldsymbol{\gamma}\,.$ (1)
[…]
@@
With a good balance of capacity and efficiency, Scaled dot-product (SDP) attention has become the most foundational attention mechanism in modern Transformers Vaswani et al., (2017). It has been widely explored in previous works and well optimized in deep learning libraries.
@@
Figure 2: Illustration for comparing different attention mechanisms: (b) visualization of attention scores. SDP attention is biased towards larger-magnitude $k_{2}$. Cos attention disregards the magnitude information. $L_{2}$ attention strikes a balance among SDP and Cos attention to attend to $k_{1}$, which has the lowest $L_{2}$ distance to the query $q$.
@@
    Softmax saturation: large token magnitudes (i.e., $\|\mathbf{q}_{i}\|$, $\|\mathbf{k}_{j}\|$) lead to large pre-softmax-logit values $\hat{\alpha}_{ij}$ and consequently extremely small gradients in backpropagation Vaswani et al., (2017).
@@
To mitigate the first and third drawbacks, existing plain Transformers heavily rely on token-wise normalization to regulate the token magnitudes. Removing LN/RMSN and/or using BatchNorm(BN) Ioffe and Szegedy, (2015) instead can induce training instability and divergence Touvron et al., 2021b ; Yao et al., (2021).
@@ 3 Method
As established in Section 2, plain Transformers exhibit suboptimal performance in graph learning, due to the architectural limitations that hinder preserving and modeling token magnitude information — a critical aspect of capturing graph structures. Consequently, previous GTs attempt to address it by adopting BN and overly sophisticated attention designs (e.g., MPNNs and/or MLP-based attention) Kreuzer et al., (2021); Rampášek et al., (2022); Ma et al., (2023). They have thus strayed away from plain Transformers, impeding the transfer of previously explored training advances and the potential unification of other foundation models.
In this work, we propose two minimal and easy-to-adapt modifications to directly address the aforementioned limitations in plain Transformers. Furthermore, we introduce an extra enhancement in the PE stem to boost the information extraction of PE. These enhancements enable plain Transformers to achieve outstanding expressive power for graph learning.
@@ 3.1 AdaRMSN
Inspired by adaptive normalization layers Dumoulin et al., (2017); de Vries et al., (2017); Peebles and Xie, (2023), we propose adaptive RMSN (AdaRMSN)
$\text{AdaRMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\cdot\gamma^{\prime}(\mathbf{x})\,,\quad\text{where }\gamma^{\prime}(\mathbf{x}):=\frac{1}{\sqrt{D}}\|\boldsymbol{\alpha}\cdot\mathbf{x}+\boldsymbol{\beta}\|\,.$ (3)
The parameter $\boldsymbol{\beta}\in\mathbb{R}^{D}$ is initialized as $\mathbf{1}$ and $\boldsymbol{\alpha}\in\mathbb{R}^{D}$ as $\mathbf{0}$, leading to $\gamma^{\prime}(\mathbf{x})=1$. AdaRMSN behaves the same as regular RMSN at the initial stage of training, but is capable of recovering the identity transformation with $\boldsymbol{\beta}=\mathbf{0}$ and $\boldsymbol{\alpha}=\mathbf{1}$ when necessary.
@@ 3.2 Simplified $L_{2}$ Attention Mechanisms
To empower SDP attention to sense both angle- and magnitude-information among query and key tokens, we revisit $L_{2}$ attention Kim et al., (2021), which is based on Euclidean distance and is capable of measuring token closeness by balancing both angles and magnitudes, in contrast to SDP attention and cosine-similarity attention (as shown in Fig. 2).
@@
This can be easily supported by existing SDP attention implementations in most deep learning libraries, e.g., PyTorch Paszke et al., (2019) 111e.g., the bias term can be directly formulated as a float attention mask in SDP-attention in PyTorch..
@@
where $\phi:\mathbb{R}^{D}\to\mathbb{R}$ and $\theta:\mathbb{R}^{D}\to\mathbb{R}$ are linear transforms. The $\phi(\mathbf{\mathbf{p}_{ij}})$ is an optional term purely based on the relative position to guarantee the universality of attention with relative PE (URPE) Luo et al., (2022). This term demands slight customization of the existing attention implementation, but we retain it since it is beneficial for learning objects with complicated structures Luo et al., (2022); Zhang et al., 2023b .
Note that, the attention with PE in relative-form is general, since it is widely employed in many existing Transformers, from language Shaw et al., (2018); Raffel et al., (2020); Press et al., (2022) to vision Dosovitskiy et al., (2021); Liu et al., (2021, 2022). Many absolute PEs are de facto explictly/implictly transformed into the relative-form in attention mechanisms Su et al., (2024); Huang et al., (2024); Zhang et al., 2024b .
@@ 3.3 Powerful Plain Graph Transformers
We follow the plain Vision Transformers architecture Dosovitskiy et al., (2021); Touvron et al., 2021a , — stems, backbone and prediction head (as shown in Fig. 1(a)(a) and Fig. 1(a)(d)).
@@
In this work, we utilize relative random walk probabilities (RRWP) Ma et al., (2023) as our demonstrating example of graph PE, considering its simplicity and effectiveness. RRWP is defined as
$\mathbf{p}^{\prime}_{ij}=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\dots,\mathbf{M}^{K-1}]_{[i,j]}\in\mathbb{R}^{K}\,,$ (6)
where $\mathbf{X}_{[i,j]}$ stands for the $i,j$th element/slice of a tensor $\mathbf{X}$; $\mathbf{M}:=\mathbf{D}^{-1}\mathbf{A}$ is the random walk matrix given the adjacency matrix $\mathbf{A}$ of the graph; and $\mathbf{I}\in\mathbb{R}^{N\times N}$ denotes the identity matrix. Most other graph PEs are applicable in our framework with minor modifications to the PE stem.
@@
Following the latest plain Transformers, we utilize the pre-norm Xiong et al., (2020) architecture, for $l=1,\cdots,L$,
@@
Before the Transformer backbone, we use small networks, usually referred to as stems, to process positional encoding $\mathbf{p}^{\prime}_{ij}$ and merge node/edge attributes $\mathbf{x}^{\prime}_{i}$/$\mathbf{e}^{\prime}_{ij}$.
@@
where $\mathtt{FC}$ stands for the fully-connected layer (e.g., linear projection); $\circ$ stands for function composition; $\mathbf{x}^{\prime}_{i}$ is treated as zero if there are no node attributes; $\mathbf{e}^{\prime}_{ij}$ is set to zero if there are no edge attributes or if $(i,j)$ is not an observed edge. Driven by the analysis of RRWP in Ma et al., (2023), we introduce additional FFNs and a final normalization layer in the PE stem to better extract the structural information, mimicking the pre-norm architecture of the Transformer backbone.
@@
We use a task-specific MLP prediction head, following GRIT Ma et al., (2023) and GraphGPS Rampášek et al., (2022): sum/mean-pooling followed by a MLP for graph level tasks; and a MLP shared by all nodes for node level tasks.
3.4 Sinusoidal PE Enhancement (SPE)

### v3

@@ Abstract
Transformers have attained outstanding performance across various modalities, owing to their simple but powerful scaled-dot-product (SDP) attention mechanisms. Researchers have attempted to migrate Transformers to graph learning, but most advanced Graph Transformers (GTs) have strayed far from plain Transformers, exhibiting major architectural differences either by integrating message-passing or incorporating sophisticated attention mechanisms. These divergences hinder the easy adoption of training advances for Transformers developed in other domains. Contrary to previous GTs, this work demonstrates that the plain Transformer architecture can be a powerful graph learner. To achieve this, we propose to incorporate three simple, minimal, and easy-to-implement modifications to the plain Transformer architecture to construct our Powerful Plain Graph Transformers (PPGT): (1) simplified $L_{2}$ attention for measuring the magnitude closeness among tokens; (2) adaptive root-mean-square normalization to preserve token magnitude information; and (3) a simple MLP-based stem for graph positional encoding. Consistent with its theoretical expressivity, PPGT demonstrates noteworthy realized expressivity on the empirical graph expressivity benchmark, comparing favorably to more complicated alternatives such as subgraph GNNs and higher-order GNNs. Its empirical performance across various graph datasets also justifies the effectiveness of PPGT. This finding underscores the versatility of plain Transformer architectures and highlights their strong potential as a unified backbone for multimodal learning across language, vision, and graph domains.
@@ 1 Introduction
Transformers have achieved excellent performance across various domains, from language Vaswani et al. (2017); Devlin et al. (2019); Brown et al. (2020) to vision Dosovitskiy et al. (2021); Touvron et al. (2021a; b), and are well known for their reduced dependency on inductive bias as well as stronger flexibility and generalizability Dosovitskiy et al. (2021). The recent success of transformer-based, multi-modal, large language models (LLMs) OpenAI (2024); Dubey et al. (2024) has also brought the once-unattainable dream of building artificial general intelligence closer to reality. However, these modalities are primarily data on Euclidean spaces, while non-Euclidean domains are still under-explored. Non-Euclidean domains like graphs are important for describing physical systems with different symmetry properties, from protein structure prediction Jumper et al. (2021) to complex physics simulation Sanchez-Gonzalez et al. (2018). Therefore, many researchers have attempted to migrate Transformer architectures for graph learning for the future integration into multimodal foundation models.
Nevertheless, unlike other Euclidean domains, the characteristics of graphs make the naive migration challenging. Following the suboptimal performance of an early attempt Dwivedi and Bresson (2021), researchers introduced significant architectural modifications to make Transformers perform well for graph learning. These include: implicit/explicit message-passing mechanisms Kreuzer et al. (2021); Rampášek et al. (2022), edge-updating Kim et al. (2022); Hussain et al. (2022); Ma et al. (2023), and sophisticated non-SDP (scaled-dot-product) attention designs Chen et al. (2022); Ma et al. (2023). These complexities hinder the integration of advances for Transformers arising from other domains and impede progress towards potential future unification of multi-modalities.
In this work, we step back and rethink the difficulties that a plain Transformer architecture faces when processing graph data. Instead of adding significant architectural modifications, we propose three minimalist but effective modifications to empower plain Transformers with advanced capabilities for learning on graphs (as shown in Fig. 1). This results in our proposed Powerful Plain Graph Transformers (PPGTs), which have remarkable empirical and theoretical expressivity for distinguishing graph structures and superior empirical performance on real-world graph datasets. These modifications can be incorporated without significantly changing the architecture of the plain Transformer. Thus, our proposal both retains the simplicity and generality of the Transformer and offers the potential of facilitating cross-modality unification.
@@ 2.1 Graph Learning
Graphs are non-Euclidean geometric spaces with irregular structures and symmetry to permutation (i.e., invariant/equivariant). A key factor for learning graphs is the ability to distinguish the distinct structures of the input graphs, which is typically referred to as the expressivity of the graph model Xu et al. (2019); Zhang et al. (2023b). Currently, most graph neural networks (GNNs) are developed based on the framework of the Weisfeiler-Leman (WL) Isomorphism test Weisfeiler and Leman (1968) – a color-refinement algorithm based on multisets $\{\!\{\cdot\}\!\}$ encoding. For example, message-passing networks (MPNNs) based on 1-WL Xu et al. (2019), distinguish graphs by encoding the neighborhood of each node as a multiset Xu et al. (2019). To go beyond 1-WL, researchers have extended to $K$-WL GNNs Morris et al. (2019), $K$-Folklore-WL (FWL) GNNs Feng et al. (2023), and Generalized-distance-WL (GD-WL) GNNs Zhang et al. (2023b), which are also based on such multiset encoding schemes, albeit with different multiset objects. It is worth mentioning that, as discussed in Xu et al. (2019); Zhang et al. (2023b), the cardinalities of multisets are crucial for distinguishing multisets, e.g., $\{\!\{a,b\}\!\}$ versus $\{\!\{a,a,b,b\}\!\}$. Generally, the cardinality is encoded into the token representation as its magnitude, e.g., encoding $\{\!\{a,b\}\!\}$ and $\{\!\{a,a,b,b\}\!\}$ as $\mathbf{x}$ and $c\cdot\mathbf{x}$ respectively, for $c\neq 1\in\mathbb{R}^{+}$. Thus, the loss of token magnitude information weakens the ability to distinguish multisets and, consequently, graph structures.
@@
The philosophy of Graph Transformers (GTs) is to learn graph representations using positional encodings (PEs) rather than operating directly on the original input graphs. Zhang et al. (2023b) provide a theoretical framework (GD-WL) that demonstrates that GTs’ stronger expressivity stems from finer-grained information beyond 1-WL encoded in graph PEs. Zhang et al. (2024b) demonstrate that GD-WL not only applies to distance-like or relative graph PEs but also enables analysis of absolute PEs (e.g., Laplacian PE) by transforming them into their relative counterparts. Concurrently, Ma et al. (2024) interpret graph PE as pseudo-coordinates in graph spaces that mimic manifolds, demonstrating that it can extend beyond graph distance and is not restricted to the Transformer architecture (e.g., convolutions on pseudo-coordinates). We follow the same philosophy and show that with a simple stem on pseudo-coordinates, even plain Transformers can achieve strong expressive power in graph learning.
@@ 2.2 Limitations in Plain Transformer Architectures
LayerNorm (LN Ba et al. (2016)) and Root-Mean-Square-Norm (RMSN Zhang and Sennrich (2019)) are two widely used token-wise normalization techniques in Transformer-based models that effectively control token magnitudes:
$\text{LN}(\mathbf{x})=\frac{\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}-\frac{1}{D}\mathbf{1}^{\intercal}\mathbf{x}\|}\odot\bm{\gamma}+\bm{\beta}\,$$\quad$ $\text{RMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\odot\bm{\gamma}\,.$ (1)
Here $\mathbf{x}\in\mathbb{R}^{D}$ are token vectors; $\|\cdot\|\in[0,\infty)$ is the $L_{2}$-norm (i.e., magnitude) of a vector; and $\bm{\gamma},\bm{\beta}\in\mathbb{R}^{D}$ are parameters of a learnable affine transform.
They retract token representations onto a hypersphere, a property essential for dot-product attention mechanisms, echoing the notion of retraction as used in computational physics. However, both LN and RMSN are strictly invariant to changes in input magnitude (see Proposition E.3 and the case study in Appx. B.2), which can result in the loss of valuable token magnitude information
@@
With a good balance of capacity and efficiency, Scaled dot-product (SDP) attention has become the most common attention mechanism in modern Transformers Vaswani et al. (2017). It has been widely explored in previous works and well optimized in deep learning libraries.
@@
Figure 2: Illustration for comparing different attention mechanisms: (b) visualization of attention scores. SDP attention is biased towards larger-magnitude $k_{2}$. Cos attention disregards the magnitude information. $L_{2}$ attention strikes a balance between SDP and Cos attention to attend to $k_{1}$, which has the lowest $L_{2}$ distance to the query $q$.
@@
    Softmax saturation: large token magnitudes (i.e., $\|\mathbf{q}_{i}\|$, $\|\mathbf{k}_{j}\|$) lead to large pre-softmax-logit values $\hat{\alpha}_{ij}$ and consequently extremely small gradients in backpropagation Vaswani et al. (2017).
@@
To mitigate the first and third drawbacks, existing plain Transformers heavily rely on token-wise normalization to regulate the token magnitudes. Removing LN/RMSN and/or using BatchNorm(BN) Ioffe and Szegedy (2015) instead can induce training instability and divergence Touvron et al. (2021b); Yao et al. (2021).
@@ 3 Method
As established in Section 2, plain Transformers exhibit suboptimal performance in graph learning, due to the architectural limitations that hinder preserving and modeling token magnitude information — a critical aspect of capturing graph structures. Consequently, previous GTs attempt to address it by adopting BN and overly sophisticated attention designs (e.g., MPNNs and/or MLP-based attention) Kreuzer et al. (2021); Rampášek et al. (2022); Ma et al. (2023). They have thus strayed away from plain Transformers, impeding the transfer of previously explored training advances and the potential unification of other foundation models.
In this work, we propose two minimal and easy-to-adapt modifications to directly address the aforementioned limitations in plain Transformers. Furthermore, we introduce an extra enhancement in the PE stem to boost the information extraction of PE. These enhancements enable plain Transformers to achieve stronger expressive power for graph learning.
@@ 3.1 AdaRMSN
Inspired by adaptive normalization layers Dumoulin et al. (2017); de Vries et al. (2017); Peebles and Xie (2023), we propose adaptive RMSN (AdaRMSN)
$\text{AdaRMSN}(\mathbf{x})=\frac{\mathbf{x}}{\frac{1}{\sqrt{D}}\|\mathbf{x}\|}\cdot\gamma^{\prime}(\mathbf{x})\,,\quad\text{where }\gamma^{\prime}(\mathbf{x}):=\frac{1}{\sqrt{D}}\|\bm{\alpha}\cdot\mathbf{x}+\bm{\beta}\|\,.$ (3)
The parameter $\bm{\beta}\in\mathbb{R}^{D}$ is initialized as $\mathbf{1}$ and $\bm{\alpha}\in\mathbb{R}^{D}$ as $\mathbf{0}$, leading to $\gamma^{\prime}(\mathbf{x})=1$. AdaRMSN behaves the same as regular RMSN at the initial stage of training, but is capable of recovering the identity transformation with $\bm{\beta}=\mathbf{0}$ and $\bm{\alpha}=\mathbf{1}$ when necessary.
@@ 3.2 Simplified $L_{2}$ Attention Mechanisms
To empower SDP attention to sense both angle- and magnitude-information among query and key tokens, we revisit $L_{2}$ attention Kim et al. (2021), which is based on Euclidean distance and is capable of measuring token closeness by balancing both angles and magnitudes, in contrast to SDP attention and cosine-similarity attention (as shown in Fig. 2).
@@
This can be easily supported by existing SDP attention implementations in most deep learning libraries, e.g., PyTorch Paszke et al. (2019) 111e.g., the bias term can be directly formulated as a float attention mask in SDP-attention in PyTorch..
@@
where $\phi:\mathbb{R}^{D}\to\mathbb{R}$ and $\theta:\mathbb{R}^{D}\to\mathbb{R}$ are linear transforms. The $\phi(\mathbf{\mathbf{p}_{ij}})$ is an optional term purely based on the relative position to guarantee the universality of attention with relative PE (URPE) Luo et al. (2022). This term demands slight customization of the existing attention implementation, but we retain it since it is beneficial for learning objects with complicated structures Luo et al. (2022); Zhang et al. (2023b). Notably, this form can be viewed as Continuous Kernel Graph Convolution Ma et al. (2024) with a dynamic density function.
Note that, the attention with PE in relative-form is general, since it is widely employed in many existing Transformers, from language Shaw et al. (2018); Raffel et al. (2020); Press et al. (2022) to vision Dosovitskiy et al. (2021); Liu et al. (2021; 2022). Many absolute PEs are de facto explicitly/implicitly transformed into the relative-form in attention mechanisms Su et al. (2024); Huang et al. (2024); Zhang et al. (2024b).
@@ 3.3 Powerful Plain Graph Transformers
We follow the plain Vision Transformers architecture Dosovitskiy et al. (2021); Touvron et al. (2021a), — stems, backbone and prediction head (as shown in Fig. 1(a) and Fig. 1(d)).
@@
In this work, we utilize relative random walk probabilities (RRWP) Ma et al. (2023) as our demonstrating example of graph PE, considering its simplicity and effectiveness. RRWP is defined as
$\mathbf{p}^{\prime}_{ij}=[\mathbf{I},{\textnormal{w}},{\textnormal{w}}^{2},\dots,{\textnormal{w}}^{K-1}]_{[i,j]}\in\mathbb{R}^{K}\,$, (6)
where $\mathbf{X}_{[i,j]}$ stands for the $i,j$th element/slice of a tensor $\mathbf{X}$; ${\textnormal{w}}:=\mathbf{D}^{-1}\mathbf{A}$ is the random walk matrix given the adjacency matrix $\mathbf{A}$ of the graph; and $\mathbf{I}\in\mathbb{R}^{N\times N}$ denotes the identity matrix. Most other graph PEs are applicable in our framework with minor modifications to the PE stem.
@@
Following the latest plain Transformers, we utilize the pre-norm Xiong et al. (2020) architecture, for $l=1,\cdots,L$,
@@
Before the Transformer backbone, we use small networks, usually referred to as stems, to process positional encoding $\mathbf{p}^{\prime}_{ij}$ and merge node/edge attributes $\mathbf{x}^{\prime}_{i}$$\mathbf{e}^{\prime}_{ij}$.
@@
where $\mathtt{FC}$ stands for the fully-connected layer (e.g., linear projection); $\circ$ stands for function composition; $\mathbf{x}^{\prime}_{i}$ is treated as zero if there are no node attributes; $\mathbf{e}^{\prime}_{ij}$ is set to zero if there are no edge attributes or if $(i,j)$ is not an observed edge. Driven by the analysis of RRWP in Ma et al. (2023), we introduce additional FFNs and a final normalization layer in the PE stem to better extract the structural information, mimicking the pre-norm architecture of the Transformer backbone.
@@
Unless otherwise specified, we employ a task-specific MLP prediction head, following the designs of GRIT Ma et al. (2023) and GraphGPS Rampášek et al. (2022) – for graph-level tasks, we apply sum or mean pooling followed by an MLP; for node-level tasks, we use an MLP shared across all nodes. For OGBN-ArXiv, we utilize a class-attention prediction head inspired by CaiT Touvron et al. (2021b), which is better suited for the graph-sampling strategy.
3.4 Anti-Spectral-Bias with Sinusoidal PE Enhancement (SPE)